# Low-Rank Expectile Representations of a Data Matrix, with Application to Diurnal Heart Rates

### v3

@@ 3.2 Accuracy of the fitted values
3.3 Comparison of Optimization Algorithms
In our pursuit to minimize the loss function, we have selected three candidate gradient-based algorithms: Broyden–Fletcher–Goldfarb–Shanno algorithm (BFGS) [Gol70], Limited-Memory Broyden–Fletcher– Goldfarb–Shanno algorithm (LBFGS) [LN89], and Conjugate Gradient method (CG) [HS+52]. To evaluate the performance and efficiency of these algorithms, we repetitively generated different simulated datasets as described above 100 times, and each time we conducted 50 initialization with specific $\tau=0.2$ and rank $k=3$. We recorded the average loss and time of each algorithm and find the algorithm with the smallest average loss and time respectively. Subsequently, we .
The aggregated results, depicted in the table 1 below, for each of the 100 different simulated datasets, the three algorithms have the minimum final loss almost evenly, and the differences in loss values among the three algorithms are extremely small, less than 0.00000001. While LBFGS outperforms significantly in the time of execution. It’s about 100 times faster than BFGS and 2-3 times faster than CG.
@@ 3.3 Comparison of Optimization Algorithms
BFGS 36 0
LBFGS 31 100
CG 33 0
Table 1: Algorithm comparison with 100 simulated datasets and 50 initialization of each dataset
@@
3.4 Resilience of Model to Random Initialization
@@
Figure 1: Heatmap of difference of final loss value and mean of absolute difference of 45 pairs
@@ 4.2.3 Intraclass Correlation Coefficient Analysis
The Intraclass Correlation Coefficient (ICC) is a statistical measure ranging from 0 to 1 that assesses the reliability or consistency of measurements made by different observers measuring the same phenomenon [Don86].
In this study, we calculated ICC values for each individual’s predictions of variables $C$ and $V$ separately to evaluate the consistency of our model’s predictions under different $\tau$ values. We constructed a new dataset by combining the prediction matrices of $C$ and $V$ with the corresponding participant IDs and detection dates. This setup allowed us to treat $Y$ as either $C$ or $V$ and $G$ as each individual participant, facilitating the computation of ICC values for different $\tau$ levels. The graph in 7 displays the ICC values with respect to different expectile values.
Figure 7: ICC Analysis Result
The ICC values for the predicted $C$ at $\tau=0.1$, $0.5$, and $0.9$ were 0.918, 0.883, and 0.646, respectively. These high ICC values, especially at $\tau=0.1$ and $\tau=0.5$, indicate a strong consistency among individuals in the model’s predictions for $C$. Despite a decrease in the ICC value at $\tau=0.9$, the consistency remains at an acceptable level, suggesting that the model maintains reasonable reliability across different $\tau$ values [DK80].
In contrast, the ICC values for the predicted $V$ at $\tau=0.1$, $0.5$, and $0.9$ were significantly lower, at 0.406, 0.348, and 0.285, respectively. These results imply that the within-group consistency of the predicted $V$ is relatively low across all $\tau$ values. The decreasing trend of ICC values with increasing $\tau$ suggests a reduction in consistency, possibly due to higher internal variability or measurement error in the predictions for $V$.
The disparity between the ICC values of predicted $C$ and $V$ indicates that our model predicts $C$ with higher reliability compared to $V$. This difference may be attributed to inherent characteristics of the variables or the model’s sensitivity to them. The lower consistency in predicted $V$ highlights the need for further investigation to understand the underlying factors contributing to this variability.
The analysis of ICC for $C$ and $V$ provides insights into the reliability of the model’s predictions across individuals and conditions. A high ICC for $C$ indicates that the additive effects, representing overall individual-level deviations, are consistently captured by the model across different expectile levels $(\tau)$. This suggests that $C$ effectively reflects stable individual characteristics. Conversely, the typically lower ICC for $V$ implies that the multiplicative effects, which account for interactions with latent factors $(U)$, exhibit greater variability both within and across individuals. This variability may arise from external influences, such as behavioral or environmental factors, that introduce complexity to the observed data. The distinction between the reliability of $C$ and $V$ underscores the complementary roles of additive and multiplicative effects in explaining circadian heart rate variations, with $C$ capturing stable individual differences and $V$ highlighting dynamic interactions and higher-order patterns.
@@ Appendix B Algorithms
Algorithm 4 Standardize Matrix Components

### v4

@@ 3.2 Accuracy of the fitted values
3.3 Performance of Loss Minimization Under Different Optimization Algorithms
In our pursuit to minimize the loss function, we have selected three candidate gradient-based algorithms: Broyden–Fletcher–Goldfarb–Shanno algorithm (BFGS) [Gol70], Limited-Memory Broyden–Fletcher– Goldfarb–Shanno algorithm (LBFGS) [LN89], and Conjugate Gradient method (CG) [HS+52]. To evaluate the performance and efficiency of these algorithms, we repetitively generated different simulated datasets as described above 100 times, and each time we conducted 100 initializations with specific $\tau=0.2$ and rank $k=3$. We recorded the minimum loss and time of each algorithm and found the algorithm with the smallest minimum loss and time respectively.
The aggregated results, depicted in the table 1 below, for each of the 100 different simulated datasets, CG has the minimum final loss almost evenly, but the differences in loss values among the three algorithms are extremely small, less than $10^{-7}$. While LBFGS outperforms significantly in the time of execution. It’s about 100 times faster than BFGS and 2-3 times faster than CG.
@@ 3.3 Performance of Loss Minimization Under Different Optimization Algorithms
BFGS 0 0
LBFGS 0 100
CG 100 0
Table 1: Algorithm comparison with 100 simulated datasets and 100 initialization of each dataset
@@
Figure 1: Heatmaps of difference of final loss value of fitted values for 45 pairs
3.4 Resilience of Fitted Values Estimates Under Random Initialization
@@
Figure 2: Heatmaps of mean of absolute difference of fitted values for 45 pairs
@@ 4.2.3 Intraclass Correlation Coefficient Analysis
The Intraclass Correlation Coefficient (ICC) offers a means to assess the consistency of predictions across different participants and days [Don86]. We computed ICC values for the additive component $C$ and the multiplicative component $V$ at $\tau=0.1,0.5,$ and $0.9$, reflecting different regions of the heart rate distribution.
Figure 8: ICC Analysis Results
For $C$, the ICC values are 0.918 at $\tau=0.1$, 0.883 at $\tau=0.5$, and 0.646 at $\tau=0.9$. These higher ICCs at lower and median expectiles suggest that $C$ captures more stable individual-level features, such as overall resting or moderately active states, which vary less across participants and days. As $\tau$ increases to 0.9, the slight drop in ICC implies that the model’s estimates of $C$ are somewhat less consistent when focusing on higher heart rate tendencies, potentially reflecting behavioral fluctuations that alter the additive baseline.
In contrast, $V$ exhibits ICCs of 0.406, 0.348, and 0.285 for $\tau=0.1,0.5,$ and $0.9$, respectively. These relatively lower and decreasing values indicate that the component interacting with the latent factors is more sensitive to day-to-day or participant-to-participant variability in human behavior. Higher $\tau$ highlights the upper tail of the heart rate distribution, where episodic behaviors such as short bursts of vigorous activity lead to greater unpredictability in $V$. Consequently, the ICC for $V$ declines as $\tau$ increases, suggesting that modeling higher expectiles encounters greater individual-level divergence.
Overall, the gap between ICC values for $C$ and $V$ indicates that the additive component is more reliably estimated, reflecting relatively stable factors over time and across individuals, while the multiplicative component captures interactions susceptible to short-term or high-intensity behavioral differences. This aligns with the model’s aim to disentangle stable baseline effects ($C$) from dynamic factors ($V$) under different regions of the heart rate distribution [DK80].
@@ Appendix B Algorithms
Algorithm 4 Canonicalize Parameter Estimates